OpenStudy (anonymous):

Roots of a cubic. Show that the function f(x)=((x-a)^2) ((x-b)^2) + (x) takes on the value ((a+b)/2) for some value of x.

OpenStudy (anonymous):

i think we can use intermediate value theorem here...are u familiar with that?

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

note that \(f(a)=a\) , \(f(b)=b\) and function is continous

OpenStudy (anonymous):

I get what you mean but how does it apply to this shoulwing that this has 3 solutions and not just 1

OpenStudy (anonymous):

oops...i meant...I get what you said...but how do i show that this has 3 solutions and not just 1?

OpenStudy (anonymous):

this a polynomial from degree 4 !!

OpenStudy (anonymous):

ok that makes sense :)

OpenStudy (anonymous):

so it should have 4 solutions !!!

OpenStudy (anonymous):

the book says it has 3 solutions and we have to show that's true

OpenStudy (anonymous):

and how does that first equation take on the value (a+b)/2 for some value of x

OpenStudy (anonymous):

\(f(a)=a\) and \(f(b)=b\) and we know that \(\frac{a+b}{2}\) lies on the interval [a,b] so according to the IVT there is a number c on interval [a,b] for which\[f(c)=\frac{a+b}{2}\]

OpenStudy (anonymous):

ohh okay but does the IVT only state that in between the a and b there lies only the value of 0 as one solution? I don't understand what the other two out of the three solutions are?

OpenStudy (anonymous):

to my understanding IVT has to do nothing with roots here

OpenStudy (anonymous):

Oh wait, I'm so sorry, I wrote the question entirely wrong. It's asking to show that the equation x^3 - 15x +1 = 0 has three solutions in the interval [-4,4]. Would this still require the IVT?

OpenStudy (anonymous):

oh :) that makes sense now

OpenStudy (anonymous):

yeah we can use IVT here

OpenStudy (anonymous):

\[f(x)=x^3-15x+1\]for example\[f(-4)=-64+60+1=-3\]\[f(-2)=-8+30+1=23\]so there is a root on interval [-4,-2] u try for intervals [-2,2] and [2,4]

OpenStudy (anonymous):

i wish could give you another Medal mukushla :)

OpenStudy (anonymous):

oh tnx man

OpenStudy (anonymous):

ohhhhhhhh so just plug in and get the values

OpenStudy (anonymous):

yeah for interval [-4,-2] f(-4)=-3 and f(-2)=23 since -3<0<23 so accordind to IVT there is a 'c' on the interval such that f(c)=0 use IVT for other 2 intervals like this good luck

OpenStudy (anonymous):

ohhhhhhh thank you soooooo much

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

nice i got it :)

OpenStudy (anonymous):

thank you again :) :D

OpenStudy (anonymous):

you are welcome :)